# thanks: Corresponding author

### v3

@@ III.4.1 The gluon quartic gauge couplings and the process to be considered
In this paper, three processes $pp\to j\ell^{+}\ell^{-}\bar{\nu}\nu$, $pp\to jj\ell^{+}\ell^{-}\bar{\nu}\nu$ and $pp\to\gamma j\ell^{+}\ell^{-}\bar{\nu}\nu$ are considered together. The Feynman diagrams are shown in Fig. 9. The three processes are combined together because they have similar backgrounds. The presence of two (anti-)neutrinos in the final state inevitably entails a certain degree of information loss, which offers an opportunity to assess the efficacy of our algorithm.
Except for the SM contribution, other processes contributing to background due to detector effects are considered, including $pp\to(\gamma)\ell\ell+jets$ with missing jets, $pp\to t\bar{t}\to jjW^{+}W^{-}\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ with b-jets mis-tagged, and $pp\to j\ell\ell\ell\nu$ with a missing charged lepton. They are listed in Table 3, with the cross-sections obtained by using MadGraph5 toolkit [78], except for the case of $t\bar{t}+jets$, in which we use inclusive cross-section of $t\bar{t}+jets$ production [127, 128, 129].
III.4.2 Training of the ansatz
@@
The training dateset is consist of events after $N_{part}$ cut. The number of events picked to form the training dataset is listed in Table 5. Training dataset has $N=65536$ vectors, with three qubits to encode one vector, with $d=8$, the number of qubits to store $\Psi_{\rm train}$ (or the ansatz $\Psi(\alpha_{i})$) is $n_{q}=19$. In this subsection, the ‘shifted-circular-alternating’ (SCA) entanglement layer [130] is used. The circuit of the ansatz is the one shown in Fig. 1. (d).
Without the amplitude amplification, the cc for an ansatz with SCA entanglement layers is about $\left((2+3l)n_{q}+2(2d-\log_{2}(d)-2)\right)c^{\prime}d$ for a test event. In this subsection, the calculation is performed by using QuEST, taking advantages of the controlled collapse operation of a simulator, we use a fixed $c^{\prime}=100$ for simplicity. If the cc of VQSN is to be less than that of its KNN counterpart (which is $d^{\prime}N=8.5\times 10^{5}$ with $d^{\prime}=13$ and $N=65536$), the ansatz should satisfy $l<17.6$. In this work, we consider three cases, $l=5$, $10$, and $15$. As a comparison, we consider (t-)KNN with $k=1$ , $10$, $100$, and $1000$.
III.4.3 Numerical results
@@
After the $N_{part}$ cut, the ansatz is trained. The fidelities for $l=5$, $10$ and $15$ ansatz are $47.12\%$, $47.47\%$ and $47.53\%$, respectively. A validation set is made which is consist of events after $N_{part}$ cut with $73400$ and $108000$ NP and SM events (all different from the training set), respectively. The normalized distributions of $\hat{a}$ and $\hat{b}$ for the validation set are shown in Fig. 10. It can be seen that, for both VQSN and t-KNN event selection strategies, $\hat{a}$ and $\hat{b}$ can be used for discriminate the signal events from background events. The distributions of $\hat{a}$ for the VQSN shows that the difference from different $l$s is small. Furthermore, the $\hat{a}$ of a background event is more closely aligned with 00 when using an ansatz.
@@
$\begin{split}\sigma(f_{0})&=\left(\sum_{p}\sigma^{(p)}_{\rm SM}+\sigma_{t\bar{t}}\varepsilon_{b}^{2}Br_{W\to\ell\nu}^{2}\varepsilon^{t\bar{t}}_{N_{part}}\varepsilon_{t\bar{t}}+\tilde{\sigma}_{bg}\right)\\ &+f_{0}\sum_{p}\sigma^{(p)}_{\rm int}+f_{0}^{2}\sum_{p}\sigma^{(p)}_{\rm NP},\end{split}$ (7)
where the summation $\sum_{p}$ is over the three processes $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$, $pp\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$, $\sigma_{t\bar{t}}$ is the inclusive cross-section of the production of $t\bar{t}+jets$ which is taken as $832\;({\rm pb})$, $\varepsilon_{b}=1-77\%$ is from mis-tagging of $b$-jets, $Br_{W\to\ell\nu}=21.7\%$ [88], $\varepsilon^{t\bar{t}}_{N_{part}}$ is the cut efficiency of $N_{part}$ cut in Table 6, $\varepsilon_{t\bar{t}}$ is the cut efficiency of (V)QSN and (t-)KNN estimated using the $t\bar{t}$ events, and $\tilde{\sigma}_{bg}$ is the sum of $\ell\ell+jets$, $j\ell\ell\ell\nu$, and $\gamma\ell\ell+jets$ backgrounds after all cuts. Then, $\sigma_{bg}=\sigma(f_{0}=0)$ and $\sigma_{s}=\sigma(f_{0})-\sigma_{bg}$,
@@
t-KNN $10$ $[-0.056,0.055]$ $[-0.069,0.067]$ $[-0.090,0.088]$
$100$ $[-0.039,0.038]$ $[-0.048,0.047]$ $[-0.064,0.063]$
$1000$ $[-0.031,0.031]$ $[-0.039,0.039]$ $[-0.053,0.053]$
Table 8: The expected constraints on the $f_{0}\;({\rm TeV}^{-4})$ at $\sqrt{s}=13\;{\rm TeV}$, $L=137\;{\rm fb}^{-1}$ and $S_{stat}=2$, $3$, and $5$ using different event selection strategies.
@@
Figure 12: Comparison of the constraints on $M_{0}$ when $f_{0}$ is positive at $\mathcal{S}_{stat}=2$ obtained from different event selection strategies. The result of Ref. [91] is shown as the dashed line.
Using the upper bound of coefficient constraint, one can set a constraint on $M_{0}$ when $f_{0}$ is positive. The results at $\mathcal{S}_{stat}=2$ are shown in Fig. 12, which provides a more clear comparison of different algorithms. It can be seen that t-KNN ($k=1000$) is the most effective, followed by t-KNN ($k=100$), VQSN ($l=15$), and VQSN ($l=10$). Furthermore, the result presented in Ref. [91] ($M_{0}\geq 1040\;{\rm MeV}$) are shown as a dashed line, and the four aforementioned cases demonstrate tighter constraints. This suggests that the processes under consideration are sensitive to the gQGCs and are worth of consideration in the global fit of the constraints on gQGCs. Furthermore, the results demonstrate the efficacy of the t-KNN and VQSN algorithms employed. The validity is a crucial consideration when using an EFT [133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144], and can usually be tested through unitarity. According to the Ref. [91], the region of coefficients in this study is free from the violation of unitarity.
@@
cc $9.9\times 10^{8}$ $8.5\times 10^{4}$ $1.6\times 10^{5}$ $2.4\times 10^{5}$ $8.5\times 10^{5}$
@@
Table 9: The number of single qubit gates, two qubit gates used in (V)QSN event selection strategies. The ‘total’ represents the number of gates to measure the classification assignment once for one test vector. The depths of circuits are also listed.
We use the average success rate of background events to estimate the cc necessary to determine the classification of one vector when $c^{\prime}=100$. The requisite number of quantum gates is presented in Table 9. When counting the number of quantum gates, the single qubit gates are combined as the single-qubit unitary gates (denoted as $U$), and the gates irrelevant of the measurements are removed. The total number of quantum gates is defined as the sum of the number of $U$ gates and the number of two-qubit gates. It can be observed that the cc for VQSN is less than that of (t-)KNN. This indicates that even with classical data, VQSN can still be more efficient than (t-)KNN. Furthermore, the cc is reduced by an order of magnitude when $l=5$. Note that, the superiority of VQSN can be even further evident for a larger training dataset or a larger $d$.
@@
III.5 Search for the anomalies quartic gauge couplings in the tri-photon process at muon colliders
III.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
As machine learning algorithms, (V)QSN and (t-)KNN should not depend on the NP or process to be searched. Thus, as an example, we consider the dimension-8 operators contributing to the anomalies quartic gauge couplings (aQGCs) in the tri-photon process at muon colliders. Due to the attention paid to vector boson scattering (VBS) processes at the LHC, the aQGCs which are suitable to be studied in VBS processes have also been widely studied phenomenologically [96, 145, 100, 102, 103] and experimentally [146, 147, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 148, 149, 150].
@@ III.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
The dimension-8 operators contributing to the aQGCs are often classified as scalar/longitudinal operators $O_{S_{i}}$, mixed transverse and longitudinal operators $O_{M_{i}}$ and transverse operators $O_{T_{i}}$. It has been found that the tri-photon processes at the muon colliders are sensitive to the $O_{T_{i}}$ operators [151], they are [152, 153, 100],
$\begin{split}&O_{T,0}={\rm Tr}\left[\widehat{W}_{\mu\nu}\widehat{W}^{\mu\nu}\right]\times{\rm Tr}\left[\widehat{W}_{\alpha\beta}\widehat{W}^{\alpha\beta}\right],\\ &O_{T,1}={\rm Tr}\left[\widehat{W}_{\alpha\nu}\widehat{W}^{\mu\beta}\right]\times{\rm Tr}\left[\widehat{W}_{\mu\beta}\widehat{W}^{\alpha\nu}\right],\\ &O_{T,2}={\rm Tr}\left[\widehat{W}_{\alpha\mu}\widehat{W}^{\mu\beta}\right]\times{\rm Tr}\left[\widehat{W}_{\beta\nu}\widehat{W}^{\nu\alpha}\right],\\ &O_{T,5}={\rm Tr}\left[\widehat{W}_{\mu\nu}\widehat{W}^{\mu\nu}\right]\times B_{\alpha\beta}B^{\alpha\beta},\\ &O_{T,6}={\rm Tr}\left[\widehat{W}_{\alpha\nu}\widehat{W}^{\mu\beta}\right]\times B_{\mu\beta}B^{\alpha\nu},\\ &O_{T,7}={\rm Tr}\left[\widehat{W}_{\alpha\mu}\widehat{W}^{\mu\beta}\right]\times B_{\beta\nu}B^{\nu\alpha},\\ &O_{T,8}=B_{\mu\nu}B^{\mu\nu}\times B_{\alpha\beta}B^{\alpha\beta},\\ &O_{T,9}=B_{\alpha\mu}B^{\mu\beta}\times B_{\beta\nu}B^{\nu\alpha},\\ \end{split}$ (8)
@@ III.5.2 Training of the ansatz
Different from the previous subsections, in the generation of the events, the standard cuts are set to be as same as the default except for the $p_{\gamma}^{T}$ cut which is set to be $p_{\gamma}^{T}\geq 0.1E_{beam}$ where $E_{beam}$ is the beam energy which can suppressing the SM background and ease the pressure on numerical computation consumption. The cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$ are considered [125, 126]. To generate the training set, $2^{16}+2^{16}$ events are chosen from the SM background and the NP events. It has been shown that, the kinematic features are similar for different operators, therefore in the training set, we use only the events from $O_{T_{0}}$. It is required that the final state to have at least three photons. The observables chosen to form the feature space are listed in Table 10, where $E_{\gamma_{1,2,3}}$ and $p^{T}_{\gamma_{1,2,3}}$ are energies and transverse momenta of the hardest, second hardest and the third hardest photons. Using this feature space, $d=6$ and $d^{\prime}=4$, and with $2^{17}$ events, the ansatz takes $n_{q}=19$ qubits. In this subsection, we also use the SCA entanglement layer and with $l=10$, the circuit is the one shown in Fig. 1. (d). After training, the fidelities are $46.8\%$, $46.8\%$, $46.9\%$ and $46.9\%$ for the cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$, respectively.
III.5.3 Numerical results
Figure 14: The cross-sections $\sigma$ after cuts compared with the fitted $\sigma$ according to Eq. (9) of $f_{T_{i}}/\Lambda^{4}$ at different collision energies, $\sqrt{s}=3\;{\rm TeV}$ (row $1$), $10\;{\rm TeV}$ (row $2$), $14\;{\rm TeV}$ (row $3$), and $30\;{\rm TeV}$ (row $4$).

### v4

@@ 3.4.1 The gluon quartic gauge couplings and the process to be considered
In this paper, three processes $p\to j\ell^{+}\ell^{-}\bar{\nu}\nu, pp$$p\to jj\ell^{+}\ell^{-}\bar{\nu}\nu$ and $pp\to\gamma j\ell^{+}\ell^{-}\bar{\nu}\nu$ are considered together. The Feynman diagrams are shown in Fig. 12. The three processes are combined together because they have similar backgrounds. The presence of two (anti-)neutrinos in the final state inevitably entails a certain degree of information loss, which offers an opportunity to assess the efficacy of our algorithm.
Except for the SM contribution, other processes contributing to background due to detector effects are considered, including $pp\to(\gamma)\ell\ell+jets$ with missing jets, $pp\to t\bar{t}\to jjW^{+}W^{-}\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ with b-jets mis-tagged, and $pp\to j\ell\ell\ell\nu$ with a missing charged lepton. They are listed in Table 3, with the cross-sections obtained by using MadGraph5 toolkit Alwall:2014hca , except for the case of $t\bar{t}+jets$, in which we use inclusive cross-section of $t\bar{t}+jets$ production Czakon:2011xx ; CMS:2020grm ; ATLAS:2024aht .
3.4.2 Training of the ansatz
@@
The training dateset is consist of events after $N_{part}$ cut. The number of events picked to form the training dataset is listed in Table 5. Training dataset has $N=65536$ vectors, with three qubits to encode one vector, with $d=8$, the number of qubits to store $\Psi_{\rm train}$ (or the ansatz $\Psi(\alpha_{i})$) is $n_{q}=19$. In this subsection, the ‘shifted-circular-alternating’ (SCA) entanglement layer Sim:2019yyv is used. The circuit of the ansatz is the one shown in Fig. 4.
Without the amplitude amplification, the cc for an ansatz with SCA entanglement layers is about $\left((2+3l)n_{q}+2(2d-\log_{2}(d)-2)\right)c^{\prime}d$ for a test event. In this subsection, the calculation is performed by using QuEST, taking advantages of the controlled collapse operation of a simulator, we use a fixed $c^{\prime}=100$ for simplicity. If the cc of VQSN is to be less than that of its KNN counterpart (which is $d^{\prime}N=8.5\times 10^{5}$ with $d^{\prime}=13$ and $N=65536$), the ansatz should satisfy $l<17.6$. In this work, we consider three cases, $l=5$, $10$, and $15$. As a comparison, we consider (t-)KNN with $k=1$ , $10$, $100$, and $1000$. The improvement brought by a larger $k$ is negligible when $k$ is larger than $1000$.
3.4.3 Numerical results
@@
After the $N_{part}$ cut, the ansatz is trained. The fidelities for $l=5$, $10$ and $15$ ansatz are $12\%, 47.47$$47\%$ and $47.53\%$, respectively. A validation set is made which is consist of events after $N_{part}$ cut with $73400$ and $108000$ NP and SM events (all different from the training set), respectively. The normalized distributions of $\hat{a}$ and $\hat{b}$ for the validation set are shown in Fig. 13. It can be seen that, for both VQSN and t-KNN event selection strategies, $\hat{a}$ and $\hat{b}$ can be used for discriminate the signal events from background events. The distributions of $\hat{a}$ for the VQSN shows that the difference from different $l$s is small. Furthermore, the $\hat{a}$ of a background event is more closely aligned with 0 when using an ansatz.
@@
[…]
where the summation $\sum_{p}$ is over the three processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$, $\sigma_{t\bar{t}}$ is the inclusive cross-section of the production of $t\bar{t}+jets$ which is taken as $832\;({\rm pb})$, $\varepsilon_{b}=1-77\%$ is from mis-tagging of $b$-jets, $Br_{W\to\ell\nu}=21.7\%$ ParticleDataGroup:2020ssz , $\varepsilon^{t\bar{t}}_{N_{part}}$ is the cut efficiency of $N_{part}$ cut in Table 6, $\varepsilon_{t\bar{t}}$ is the cut efficiency of (V)QSN and (t-)KNN estimated using the $t\bar{t}$ events, and $\tilde{\sigma}_{bg}$ is the sum of $\ell\ell+jets$, $j\ell\ell\ell\nu$, and $\gamma\ell\ell+jets$ backgrounds after all cuts. Then, $\sigma_{bg}=\sigma(f_{0}=0)$ and $\sigma_{s}=\sigma(f_{0})-\sigma_{bg}$,
@@
t-KNN $10$ [0.056,0.055][-0.056,0.055] [0.069,0.067][-0.069,0.067] [0.090,0.088][-0.090,0.088]
$100$ [0.039,0.038][-0.039,0.038] [0.048,0.047][-0.048,0.047] [0.064,0.063][-0.064,0.063]
$1000$ [0.031,0.031][-0.031,0.031] [0.039,0.039][-0.039,0.039] [0.053,0.053][-0.053,0.053]
Table 8: The expected constraints on the $f_{0}\;({\rm TeV}^{-4})$ at $\sqrt{s}=13\;{\rm TeV}$, $L=137\;{\rm fb}^{-1}$ and $S_{stat}=2$, $3$, and 55 using different event selection strategies.
@@
Figure 15: Comparison of the constraints on $M_{0}$ when $f_{0}$ is positive at $\mathcal{S}_{stat}=2$ obtained from different event selection strategies. The result of Ref. Ellis:2018cos is shown as the dashed line.
Using the upper bound of coefficient constraint, one can set a constraint on $M_{0}$ when $f_{0}$ is positive. The results at $\mathcal{S}_{stat}=2$ are shown in Fig. 15, which provides a more clear comparison of different algorithms. It can be seen that t-KNN ($k=1000$) is the most effective, followed by t-KNN ($k=100$), VQSN ($l=15$), and VQSN ($l=10$). Furthermore, the result presented in Ref. Ellis:2018cos  ($M_{0}\geq 1040\;{\rm MeV}$) are shown as a dashed line, and the four aforementioned cases demonstrate tighter constraints. This suggests that the processes under consideration are sensitive to the gQGCs and are worth of consideration in the global fit of the constraints on gQGCs. Furthermore, the results demonstrate the efficacy of the t-KNN and VQSN algorithms employed. The validity is a crucial consideration when using an EFT Dong:2023nir ; Yang:2021pcf ; Guo:2020lim ; Yue:2021snv ; Fu:2021mub ; Yang:2022ilt ; Layssac:1993vfp ; Corbett:2017qgl ; Perez:2018kav ; Almeida:2020ylr ; Kilian:2018bhs ; Kilian:2021whd , and can usually be tested through unitarity. According to the Ref. Ellis:2018cos , the region of coefficients in this study is free from the violation of unitarity.
@@
$c$ $9.9\times 10^{8}$ $8.5\times 10^{4}$ $1.6\times 10^{5}$ $2.4\times 10^{5}$ $8.5\times 10^{5}$
@@
Table 9: The number of single-qubit gates, two-qubit gates used in (V)QSN event selection strategies. The ‘total’ represents the number of gates to measure the classification assignment once for one test vector. The depths of circuits are also listed.
We use the average success rate of background events to estimate the cc necessary to determine the classification of one vector when $c^{\prime}=100$. The requisite number of quantum gates is presented in Table 9. When counting the number of quantum gates, the single-qubit gates are combined as the single-qubit unitary gates (denoted as $U$), and the gates irrelevant of the measurements are removed. The total number of quantum gates is defined as the sum of the number of $U$ gates and the number of two-qubit gates. It can be observed that the cc for VQSN is less than that of (t-)KNN. This indicates that even with classical data, VQSN can still be more efficient than (t-)KNN. Furthermore, the cc is reduced by an order of magnitude when $l=5$. Note that, the superiority of VQSN can be even further evident for a larger training dataset or a larger $d$.
@@
Given the substantially higher error rate of two-qubit gates compared to single-qubit gates, evaluating the circuit depth based on two-qubit gates provides a better estimate of the overall error. For $d$-dimensional normalized complex vector (2d12d-1 dimensional feature space), $n_{q}=\log_{2}(Md)$. For pairwise entanglement layer, $dep_{ans}=l$ in the sense of two-qubit gates, and $dep_{ans}=n_{q}l$ for the case of SCA entanglement layer. $dep_{ae}=2\left(d-1-\log_{2}(d)\right)$ in the sense of two-qubit gates. The number of qubits to be measured is $\log_{2}d+1$ $l$ depends on the expressive power of the ansatz and the distribution of the training data. If we assume that the ansatz can express the training dataset with a moderate $l$, then both the error and computational complexity scale logarithmically with $M$. This is the reason why the ansatz leads to optimization.
3.5 Search for the anomalies quartic gauge couplings in the tri-photon process at muon colliders
3.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
As machine learning algorithms, (V)QSN and (t-)KNN should not depend on the NP or process to be searched. Thus, as an example, we consider the dimension-8 operators contributing to the anomalies quartic gauge couplings (aQGCs) in the tri-photon process at muon colliders. Due to the attention paid to vector boson scattering (VBS) processes at the LHC, the aQGCs which are suitable to be studied in VBS processes have also been widely studied phenomenologically Green:2016trm ; Chang:2013aya ; Anders:2018oin ; Zhang:2018shp ; Bi:2019phv and experimentally ATLAS:2014jzl ; CMS:2020gfh ; ATLAS:2017vqm ; CMS:2017rin ; CMS:2020ioi ; CMS:2016gct ; CMS:2017zmo ; CMS:2018ccg ; ATLAS:2018mxa ; CMS:2019uys ; CMS:2016rtz ; CMS:2017fhs ; CMS:2019qfk ; CMS:2020ypo ; CMS:2020fqz .
@@ 3.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
The dimension-8 operators contributing to the aQGCs are often classified as scalar/longitudinal operators $O_{S_{i}}$, mixed transverse and longitudinal operators $O_{M_{i}}$ and transverse operators $O_{T_{i}}$. It has been found that the tri-photon processes at the muon colliders are sensitive to the $O_{T_{i}}$ operators Yang:2020rjt , they are Eboli:2006wa ; Eboli:2016kko ; Anders:2018oin ,
[…]
@@ 3.5.2 Training of the ansatz
Different from the previous subsections, in the generation of the events, the standard cuts are set to be as same as the default except for the $p_{\gamma}^{T}$ cut which is set to be $p_{\gamma}^{T}\geq 0.1E_{beam}$ where $E_{beam}$ is the beam energy which can suppressing the SM background and ease the pressure on numerical computation consumption. The cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$ are considered Black:2022cth ; Accettura:2023ked . To generate the training set, $2^{16}+2^{16}$ events are chosen from the SM background and the NP events. It has been shown that, the kinematic features are similar for different operators, therefore in the training set, we use only the events from $O_{T_{0}}$. It is required that the final state to have at least three photons. The observables chosen to form the feature space are listed in Table 10, where $E_{\gamma_{1,2,3}}$ and $p^{T}_{\gamma_{1,2,3}}$ are energies and transverse momenta of the hardest, second hardest and the third hardest photons. Using this feature space, $d=6$ and $d^{\prime}=4$, and with $2^{17}$ events, the ansatz takes $n_{q}=19$ qubits. In this subsection, we also use the SCA entanglement layer and with $l=10$, the circuit is the one shown in Fig. 4. After training, the fidelities are $\%, 46.846.8\%$, $46.9\%$ and $46.9\%$ for the cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$, respectively.
3.5.3 Numerical results
Figure 17: The cross-sections $\sigma$ after cuts compared with the fitted $\sigma$ according to Eq. (9) of $f_{T_{i}}/\Lambda^{4}$ at different collision energies, $\sqrt{s}=3\;{\rm TeV}$ (row 11), $10\;{\rm TeV}$ (row 22), $14\;{\rm TeV}$ (row 33), and $30\;{\rm TeV}$ (row 44).